# In-Context Translation: Towards Unifying Image Recognition, Processing, and Generation

### v1

@@ 4.2 Results and Analyses
Visual Recognition Results. We assess the proposed ICT on three visual recognition tasks described in Sec. 3.1. Standard metrics are adopted for evaluation: (1) mean Intersection-over-Union (mIoU) for semantic segmentation; (2) root mean squared error (RMSE), absolute relative error (REL), and the accuracy under the threshold ($\delta_{1}<1.25$) for depth estimation. Quantitative and qualitative comparison results are presented in Tab. 2 and Fig. 3(a) respectively. We make the following Observations. O1: ICT outperforms PromptDiffusion by a large margin, despite both adopting the pre-trained SD as the backbone. As shown in Fig. 3(a), PromptDiffusion tends to mispredict pixels as “black”. We attribute ICT’s superior performance to our more favorable in-context image translation framework that better integrates the image context with the query image through spatial concatenation rather than the pixel-wise sum operation applied in PromptDiffusion (as we discussed in Sec. 2).. O2: The official results of Painter are better than ours, but Painter requires training on a considerable number of machines, i.e., around 128 RTX 3090 GPUs. Hence, we retrain Painter following its official code to compare with ICT more fairly under the same academia-level compute (4 RTX 3090 GPUs). In this case, our ICT is highly comparable with Painter on visual recognition tasks. The visual results presented in Fig. 3(a) also verify that ICT succeeds in perceiving various scenes regarding semantics and salient object components, and subsequently produces accurate predictions in RGB format. O3: ICT as well as Painter and PromptDiffusion still falls behind most specialized methods. However, the primary focus of this paper is to reveal the potential of generative modeling in building a general framework for three distinct categories of vision tasks, with achieving state-of-the-art performance being of our lower priority.
@@
It is worth noting that ICT performs well on keypoint detection (see Fig. 3(a) and the supplementary). Nevertheless, generating heatmaps to calculate metrics such as average precision (AP) is difficult to accomplish with the autoencoder of pre-trained SD as it introduces lossy compression. Limited by this, we do not report quantitative results. This issue can be alleviated by resorting to better pre-trained models in the future or employing an extra model to transfer the output to heatmaps as done in [23].
@@
Low-level Image Processing Results. We exploit the ability of ICT to perform low-level image processing on three image restoration tasks. Standard metrics, including PSNR, SSIM, and LPIPS [82], are used for evaluation. Tab. 3 presents the quantitative comparison results. Similar to the observations in visual recognition, here ICT attains competitive performance compared to Painter (retrained version) and surpasses PromptDiffusion in all metrics. It is worth noting that there is an upper bound for ICT because the autoencoder of pre-trained SD brings information loss (as pointed out in [23]). We apply the autoencoder to reconstruct the ground truth and calculate the metrics as our upper bound. In addition, ICT-sc produces better results than ICT-st on image deraining and denoising, indicating the capability of ICT-sc to perform versatile low-level image processing using a single model. Visual results illustrated in Fig. 3(b) also demonstrate the efficacy of ICT in handling low-level image processing tasks.
@@
Conditional Image Generation Results. We evaluate the conditional image generation performance of ICT given various conditions, including segmentation mask, depth map, keypoint, and HED edge. The commonly used Fréchet Inception Distance (FID) [25] is adopted to assess the realism of the generated images. The comparison results are presented in Table 4 and Fig. 3(c). Observations are elaborated in the following. O1: The proposed ICT achieves exceptional performance on all tasks and even surpasses the specialized method (ControlNet) on mask/depth/pose-to-image synthesis, indicating that ICT fully unleashes the generative power of pre-trained SD. O2: Painter, which is built on top of pre-trained MAE, falls short of synthesizing realistic images from conditions with sparse semantics, resulting in poor FID values. The blurry images synthesized by Painter shown in Fig. 3(c) further validate the collapse of Painter on conditional image generation tasks. O3: ICT-sc attains a higher performance than ICT-st on most of conditional image generation tasks. This showcases the effectiveness of ICT-sc in translating flexible control signals into high-fidelity images using a unified model. As presented in Fig. 3(c), given different conditions, ICT-sc manages to recognize the underlying task and synthesize photorealistic images that spatially conform to the control signals.
Results of Multi-category Model. We collect data from depth estimation, denoising, and mask-to-image synthesis to jointly train a multi-category model ICT-mc. As shown in Tab. 5, ICT-mc achieves a competitive performance very close to ICT-st, confirming the proposed framework’s ability to automatically identify the specific task through the given context and produce the corresponding desired output. It is encouraging to see the results of ICT-mc trained for these tasks involving disparate visual signals and data domains, and we believe that unifying discrimination and generation via a single model will be made possible if the proposed ICT can be trained with sufficient computational resources.
@@
Table 5: Results of our multi-category model ICT-mc. We select three representative tasks from three categories to investigate ICT-mc.
@@
ICT-st 0.420 0.135 0.857 34.55 0.907 0.095 29.9
ICT-mc 0.421 0.131 0.863 34.58 0.909 0.095 30.4
Table 6: Ablation study results of ICT-st for semantic segmentation on ADE20K.
Method Masking strategy Type of text context
region-wise whole image no prompt task-level instance-level
@@
Table 7: Ablation study results of ICT-mc regarding text context.
@@
ICT-mc w/ text context 0.421 0.131 0.863 34.58 0.909 0.095 30.4
ICT-mc w/o text context 0.466 0.154 0.826 34.36 0.907 0.101 31.5
Ablation Study. We perform ablations on two ingredients of ICT: the masking strategy and the design of text context. Instead of masking the whole image $I_{gt}$ during training, we randomly mask out a portion of $I_{gt}$ to train ICT for the semantic segmentation task. As reported in Tab. 6, this region-wise masking results in a significant performance drop, highlighting the importance of our masking strategy in unleashing the unifying ability of pre-trained SD. We also study the utility of text context by conducting two ablations for ICT-st and ICT-mc respectively. Results are provided in Tab. 6 and Tab. 7. As can be seen, text context is beneficial for some tasks (e.g., depth estimation), but the gain by applying text context is very marginal for other tasks (e.g., denoising). Therefore, one can optionally apply text context to strike a balance between high-end performance and extra human efforts.
Figure 4: Example results from our ICT in performing out-of-distribution inference. More results are given in the supplementary.
Table 8: Comparison results on out-of-distribution datasets.
@@
Table 9: Scaling behavior of ICT w.r.t. computing resources.
@@
Generalization Capability. We explore ICT’s generalization capability by applying it to unseen tasks/data. As demonstrated in Fig. 4, ICT is capable of (1) generating realistic images from the normal map that is unseen during training, (2) conducting depth estimation/keypoint detection/denoising on in-the-wild images that have different data distribution compared to the training dataset, and (3) performing segmentation/deraining on images from new domains (e.g., Van Gogh’s paintings in Fig. 4(c)). These promising results indicate that ICT learns the underlying “structure” of various visual signals and generalizes well to new scenarios by leveraging the pre-trained knowledge to conduct in-context image translation. In addition, we evaluate different methods on unseen datasets and report the quantitative results in Tab. 8. Compared to other competing methods, ICT exhibits superior performance on data unseen in training, indicating the strong generalization ability of our proposed method.
Efficiency and Scalability. Under limited compute (i.e., 4 RTX 3090 GPUs), ICT exhibits better performance (especially its unification of more categories of tasks) compared to its competitors, which showcases the higher efficiency of the proposed framework. We further conduct an experiment where ICT is trained using different amounts of computing resources. We present the results in Tab. 9. ICT shows impressive scaling behavior where the performance improves with increasing computing power.
5 Conclusion
@@
Supplementary Materials
In this supplementary material, we first provide additional implementation details of the proposed In-Context Translation (ICT) framework (Appendix 0.A), including the detailed process of the construction of context-based data for each task and training and inference details. After that, we present additional experimental results to further demonstrate the effectiveness of ICT (Appendix 0.B).
Appendix 0.A Additional Implementation Details
0.A.1 Context-based Data for Each Task
This is supplementary to Sec. 3.1 in the main paper. In the following, we provide details on how we construct context-based data for various vision tasks.
Semantic Segmentation. In this task, we transfer per-pixel semantic labels to RGB images by binding each pixel with a unique color determined by its semantic class. We utilize the semantic-color codebook released by Painter [66]. The instance-level prompt is adopted as the text context for this task. We derive the prompt following the template “semantic segmentation map of {caption}”, where the caption is obtained by applying BLIP [33] to the query image.
Depth Estimation. The depth map from NYUv2 is a one-channel image with the pixel value ranging from 00 to $10000$. To obtain an RGB image from the depth map, we linearly scale the value of each pixel to $\left[0,255\right]$ and then let R, G, B have the same re-scaled value. The text context for depth estimation is a task-level prompt: “depth map”.
Keypoint Detection. To convert keypoints into RGB images, we draw colored squares at the location of each keypoint. Each square occupies 9$\times$9 pixels and its color is determined by the semantic category of that keypoint, and we adopt the same mapping strategy used for semantic segmentation. The text context for keypoint detection is a task-level prompt: “keypoint”.
Low-level Image Processing. We use the task-level prompt as the text context for three low-level image processing tasks. “a clean image without rain”, “a sharp image without noise”, and “a bright image” are applied for image deraining, image denoising, and low-light image enhancement, respectively.
Conditional Image Generation. We adopt the same method used on visual recognition tasks to translate conditions to RGB images. Instance-level prompt, which is the caption of the output image obtained through BLIP, is used as the text context for these conditional image generation tasks.
0.A.2 Training and Inference Details
This is supplementary to Sec. 4.1 in the main paper. We adopt a smooth L1 version of Eq. (2) to train our model. During training, we randomly drop 10% text-conditioning to improve classifier-free guidance sampling [27]. We load pre-trained weights from Stable Diffusion-v1.5-inpainting for fine-tuning. We randomly sample an input-output pair as the image instruction during training and adopt a fixed pair from the training dataset (same as Painter [66]) as the image context at the inference time. By setting the random noise in the reverse diffusion process to 0 (i.e., a deterministic sampling), DDIM [59] manages to generate an image with fewer sampling steps compared to DDPM. We adopt DDIM sampling with 50 steps for inference.
Appendix 0.B Additional Results
Table 10: Results of ICT using different shots of image context.
# of shots during inference Depth estimation Denoising
[…]
one-shot 0.420 0.135 0.857 34.55 0.907 0.095
two-shot 0.416 0.131 0.863 34.73 0.911 0.095
four-shot 0.413 0.130 0.863 34.75 0.911 0.095
Additional Comparison Results. This is supplementary to Sec. 4.2 (Fig. 3 in particular) in the main paper. Here we present additional comparison results on semantic segmentation (Fig. 5), depth estimation (Fig. 6), keypoint detection (Fig. 7), low-light image enhancement (Fig. 8), image deraining (Fig. 9), image denoising (Fig. 10), mask-to-image synthesis (Fig. 11), depth-to-image synthesis (Fig. 12), pose-to-image synthesis (Fig. 13), and edge-to-image synthesis (Fig. 14). These results further demonstrate the capability of ICT to perform these three distinct categories of computer vision tasks, while the competing methods Painter [66] and PromptDiffusion [69] collapse on at least one of them.
Additional Generalization Results. This is supplementary to Fig. 4 in the main paper. Here we showcase more generalization results from ICT in Fig. 15. ICT exhibits promising performance when being applied to images from new domains/in-the-wild images and novel tasks with unseen conditions. This strong generalization capability, which is analogous to LLMs, again validates our design of building a general framework for vision tasks. Gathering more diverse data for training ICT could be promising to enhance generalizability and we plan to investigate this in future work.
Few-shot In-context Inference. Here we extend ICT to perform few-shot in-context inference. This is fulfilled by establishing the grid image with more input-output pairs. We report the results of ICT under one-shot, two-shot, and four-shot settings in Tab. 10. We observe that there is a slight gain by introducing more image context, and we think this could be further explored during training, which we leave to future work.
Text-conditional Image Generation Results. We also explore an additional task: text-conditional image generation. This task can be fulfilled by directly applying ICT-sc trained on four conditional image generation tasks where the query is set to a black image. We achieve a FID of 27.47 on the COCO validation set. This could be further improved by training a ICT-st on this task. We present some text-to-image generation results obtained from our method in Fig. 16.
Figure 5: Visual comparison results on semantic segmentation. PromptDiffusion [69] tends to mispredict pixels as “black”. Red arrow indicates several mispredictions produced by Painter [66].
Figure 6: Visual comparison results on depth estimation. The competing methods, especially PromptDiffusion [69], fail to consistently produce accurate depth predictions (red arrow).
Figure 7: Visual comparison results on keypoint detection. PromptDiffusion [69] tends to overpredict keypoints (red arrow). In contrast, the proposed ICT consistently produces accurate predictions.
Figure 8: Visual comparison results on low-light image enhancement.
Figure 9: Visual comparison results on image deraining.
Figure 10: Visual comparison results on image denoising. PromptDiffusion [69] tends to generate images with color shifts that do not adhere to the query image.
Figure 11: Visual comparison results on mask-to-image synthesis.
Figure 12: Visual comparison results on depth-to-image synthesis.
Figure 13: Visual comparison results on pose-to-image synthesis.
Figure 14: Visual comparison results on edge-to-image synthesis.
Figure 15: Additional generalization results from ICT.
Figure 16: Text-to-image generation results of ICT.

### v2

@@ IV-B Results and Analyses
Quantitative and qualitative comparison results are presented in Table II and Fig. 3 respectively. We make the following Observations. O1: ICT outperforms PromptDiffusion by a large margin, despite both adopting the pre-trained SD as the backbone. As shown in Fig. 3(a), PromptDiffusion tends to mispredict pixels as “black”. We attribute ICT’s superior performance to our more favorable in-context image translation framework that better integrates the image context with the query image through spatial concatenation rather than the pixel-wise sum operation applied in PromptDiffusion (as we discussed in Section II).. O2: The official results of Painter are better than ours, but Painter requires training on a considerable number of machines, i.e., around 128 RTX 3090 GPUs. Hence, we retrain Painter following its official code to compare with ICT more fairly under the same academia-level compute (4 RTX 3090 GPUs). In this case, our ICT is highly comparable with Painter on visual recognition tasks. The visual results presented in Fig. 3 also verify that ICT succeeds in perceiving various scenes regarding semantics and salient object components, and subsequently produces accurate predictions in RGB format. O3: ICT as well as Painter and PromptDiffusion still falls behind most specialized methods. However, the primary focus of this paper is to reveal the potential of generative modeling in building a general framework for three distinct categories of vision tasks, with achieving state-of-the-art performance being of our lower priority.
@@
It is worth noting that ICT performs well on keypoint detection (see Fig. 3(c)). Nevertheless, generating heatmaps to calculate metrics such as average precision (AP) is difficult to accomplish with the autoencoder of pre-trained SD as it introduces lossy compression. Limited by this, we do not report quantitative results. This issue can be alleviated by resorting to better pre-trained models in the future or employing an extra model to transfer the output to heatmaps as done in InstructDiffusion [38].
Low-level Image Processing Results. We exploit the ability of ICT to perform low-level image processing on three image restoration tasks. Standard metrics, including PSNR, SSIM, and LPIPS [96], are used for evaluation.
@@
Table III presents the quantitative comparison results. Similar to the observations in visual recognition, here ICT attains competitive performance compared to Painter (retrained version) and surpasses PromptDiffusion in all metrics. It is worth noting that there is an upper bound for ICT because the autoencoder of pre-trained SD brings information loss (as pointed out in InstructDiffusion [38]). We apply the autoencoder to reconstruct the ground truth and calculate the metrics as our upper bound. In addition, ICT-sc produces better results than ICT-st on image deraining and denoising, indicating the capability of ICT-sc to perform versatile low-level image processing using a single model. Visual results illustrated in Fig. 4 also demonstrate the efficacy of ICT in handling low-level image processing tasks.
Conditional Image Generation Results. We evaluate the conditional image generation performance of ICT given various conditions, including segmentation mask, depth map, keypoint, and HED edge. The commonly used Fréchet Inception Distance (FID) [97] is adopted to assess the realism of the generated images. To evaluate the alignment of the generated images with the given conditions, we follow Uni-ControlNet [98] to employ the following metrics for four different conditional image generation tasks: 1) mean Intersection-over-Union (mIoU) for mask-to-image synthesis; 2) root mean squared error (RMSE) for depth-to-image synthesis; 3) mean Average Precision (mAP) for pose-to-image synthesis; 4) SSIM for edge-to-image synthesis. We calculate these metrics by comparing the extracted conditions from the generated images and the given conditions.
@@
The comparison results are presented in Table IV and Fig. 5. Observations are elaborated in the following. O1: The proposed ICT achieves exceptional performance on all tasks and even surpasses the specialized method (ControlNet) on mask/depth/pose-to-image synthesis, indicating that ICT fully unleashes the generative power of pre-trained SD. O2: Painter, which is built on top of pre-trained MAE, falls short of synthesizing realistic images from conditions with sparse semantics, resulting in poor FID values. The blurry images synthesized by Painter shown in Fig. 5 further validate the collapse of Painter on conditional image generation tasks. O3: ICT-sc attains a higher performance than ICT-st on most of conditional image generation tasks. This showcases the effectiveness of ICT-sc in translating flexible control signals into high-fidelity images using a unified model. As presented in Fig. 5, given different conditions, ICT-sc manages to recognize the underlying task and synthesize photorealistic images that spatially conform to the control signals.
@@
TABLE V: Results of Our Multi-Category Model ICT-mc. We Select Three Representative Tasks from Three Categories to Investigate ICT-mc
@@
ICT-st 0.420 0.135 0.857 34.55 0.907 0.095 29.9 28.7
ICT-mc 0.421 0.131 0.863 34.58 0.909 0.095 30.4 33.2
Results of Multi-category Model. We collect data from depth estimation, denoising, and mask-to-image synthesis to jointly train a multi-category model ICT-mc. As shown in Table V, ICT-mc achieves a competitive performance very close to ICT-st, confirming the proposed framework’s ability to automatically identify the specific task through the given context and produce the corresponding desired output. It is encouraging to see the results of ICT-mc trained for these tasks involving disparate visual signals and data domains, and we believe that unifying discrimination and generation via a single model will be made possible if the proposed ICT can be trained with sufficient computational resources.
TABLE VI: Ablation Study Results of ICT-st for Semantic Segmentation on ADE20K
Method Masking strategy Type of text context Type of image context
region-wise whole image no prompt task-level instance-level fixed pair random pair
@@
TABLE VII: Ablation Study Results of ICT-mc Regarding Text Context
@@
ICT-mc w/ text context 0.421 0.131 0.863 34.58 0.909 0.095 30.4 33.2
ICT-mc w/o text context 0.466 0.154 0.826 34.36 0.907 0.101 31.5 29.5
IV-C Ablation Study
We perform ablations on three ingredients of ICT: the masking strategy, the design of text context, and the type of image context.
Masking Strategy. Instead of masking the whole image $I_{gt}$ during training, we randomly mask out a portion of $I_{gt}$ to train ICT for the semantic segmentation task. As reported in Table VI, this region-wise masking results in a significant performance drop, highlighting the importance of our masking strategy in unleashing the unifying ability of pre-trained SD.
Design of Text Context. We also study the utility of text context by conducting two ablations for ICT-st and ICT-mc respectively. Results are provided in Table VI and Table VII. As can be seen, text context is beneficial for some tasks (e.g., depth estimation), but the gain by applying text context is very marginal for other tasks (e.g., denoising). Therefore, one can optionally apply text context to strike a balance between high-end performance and extra human efforts.
Type of Image Context. We further investigate the effect of image context by comparing two methods of selecting image context during inference: 1) randomly selecting an input-output pair from the training data each time, and 2) adopting a fixed input-output pair from the training data for all testing samples. As shown in Table VI, applying random selection of image context attains better performance than using fixed context. This implies the potential of improving the performance of our work by finding better image context.
Figure 6: Example results from our ICT in performing out-of-distribution inference.
Figure 7: Text-to-image generation results of ICT.
TABLE VIII: Comparison Results on Out-of-Distribution Datasets
@@ IV-C Ablation Study
TABLE IX: Results of ICT Using Different Shots of Image Context
# of shots during inference Depth estimation Denoising
[…]
one-shot 0.420 0.135 0.857 34.55 0.907 0.095
two-shot 0.416 0.131 0.863 34.73 0.911 0.095
four-shot 0.413 0.130 0.863 34.75 0.911 0.095
TABLE X: Scaling Behavior of ICT w.r.t. Computing Resources
@@
IV-D Generalization Capability
We explore ICT’s generalization capability by applying it to unseen tasks/data. As demonstrated in Fig. 6, ICT is capable of 1) generating realistic images from conditions that are unseen during training (e.g., normal map, canny edge, and scribble), 2) conducting image recognition and processing (e.g., depth estimation, keypoint detection, and image denoising) on in-the-wild images that have different data distribution compared to the training dataset, and 3) performing image recognition and processing (e.g., semantic segmentation, depth estimation, and image deraining) on images from new domains (e.g., Van Gogh’s paintings in Fig. 6(c)). These promising results indicate that ICT learns the underlying “structure” of various visual signals and generalizes well to new scenarios by leveraging the pre-trained knowledge to conduct in-context image translation.
In addition, we evaluate different methods on unseen datasets and report the quantitative results in Table VIII. Compared to other competing methods, ICT exhibits superior performance on data unseen in training, indicating the strong generalization ability of our proposed method.
We also explore an additional task: text-conditional image generation. This task can be fulfilled by directly applying ICT-sc trained on four conditional image generation tasks where the query is set to a black image. We achieve a FID of 27.47 on the COCO validation set. This could be further improved by training a ICT-st on this task. We present some text-to-image generation results obtained from our method in Fig. 7.
IV-E Few-shot In-context Inference
Here we extend ICT to perform few-shot in-context inference. This is fulfilled by establishing the grid image with more input-output pairs. We report the results of ICT under one-shot, two-shot, and four-shot settings in Table IX. We observe that there is a slight gain by introducing more image context, and we think this could be further explored during training, which we leave to future work.
IV-F Efficiency and Scalability
Under limited compute (i.e., 4 RTX 3090 GPUs), ICT exhibits better performance (especially its unification of more categories of tasks) compared to its competitors, which showcases the higher efficiency of the proposed framework. We further conduct an experiment where ICT is trained using different amounts of computing resources. We present the results in Table X. ICT shows impressive scaling behavior where the performance improves with increasing computing power.
V Conclusion